<p>Biased detectors this morning to 150V, the leakage current for all detectors appears stable and is ~3.5 uA. We have also replaced the CAEN bias module so that all channels are now functional. Detectors are connected as follows;</p>
<p>CHN0 -> Top left</p>
<p>CHN1 -> Top Right</p>
<p>CHN2 -> Bottom Left</p>
<p>CHN3 -> Bottom Right</p>
<p>Reconnected the motors and moved the detectors while biased. No observable change in the leakage current during movement at 150V. An increase in the pressure of the electron target was observed during movement (up to 1e-9 mbar). This increase was only momentarily, and the pressure recovered after a couple of seconds. We expect the pressure increase from movement to decrease over time.</p>
<p> </p> |